Sep 17, 1991 - Blue Sheet

PDF Download

Project Progress Summary

September 17, 1991

As of 30 August 1991, the CEBAF Project is 60% complete and 74% obligated.

FET Operations

  • Two runs were made this week with 100 kV DC beam. Both were to search for 60 Hz modulation effects on the beam intensity and steering. One run established that the voltage modulation on the 100 kV gun is no worse than 10-4. In this run, we also found that we could deliver significantly larger beam currents by removing one of the emittance-defining apertures. We inferred average currents of over 600 microamps, which would permit sending up to 200 microamps to any one end station. It was not possible to measure the emittance, but this will be done later in the FET. In a second run, we were able to verify that we have small steering effects from 60 Hz magnetic fields at the gun. These cause a 60 Hz intensity modulation of a few percent after the emittance filter and will have to be eliminated.
  • Recirculator installation is nearing completion. All stands are in place and very few items remain to be installed. The system is over 90% under vacuum, some magnets have been operated under computer control, and most diagnostics are wired up. This week we will complete the installation and alignment and operate all of the elements from the computer.
  • The new emittance aperture in the 100 kV line should be ready for installation this week.

Software:

  • Preparations have been underway for today's TACL workshop.

WBS 1

  • Prepared 10 complete RF window assemblies for Cr2O3f Cr charge-dissipation coating. These windows will be sputtered and turned over to the cavity assembly group this week for testing.
  • Completed the order for 400 raw ceramics from Astromet for RF windows. The ceramic is 99.96% high purity alumina with 4 ppm Mg02. Completed ceramics are scheduled to arrive at CEBAF the last week of November (i.e., fired, machined, and metallized). Preparation of the clean room for window and HOM load activities is going well. Systems have been installed, leak checked and readied for final hookup. Sixty HOM flanges were machined and brazed this week. Final surf ace preparation and brazing should be complete by 1 November.
  • Ten ceramic tiles received from Ceradyne meet specifications and will be machined into load sets this week.
  • Cavity pair parts continue to be formed, machined, and welded.
  • Twelve cavities are undergoing internal inspection, pre-chemistry, and measuring for alignment in preparation for production.
  • Reworked the "Phoenix" cryomodule waveguide vacuum. Prepared all-installed cryomodule insulation vacuums for cooldown.
  • Installed new warm windows (polyethylene) on the quarter-cryomodule for vacuum and RF evaluation.
  • Six cooldown/warmup cycles completed on the quarter-cryomodule thermal cycle test.

WBS 2

Magnets:

  • Two sets of paired contracts for small magnets were awarded. The linac corrector dipole (BT) coils will be made by Millhouse Co., and Automation Service Co. will make the corresponding cores and complete the assembly. The linac quadrupole (QB) coils will be made by Millhouse Co., and United Magnet Technology will make the corresponding cores and complete the assembly.

Stands:

  • Quadrupole girders were awarded to South Technology International.
  • The first production lot of dipole stands for the west arc was received from Capital Engineering and Manufacturing.

Vacuum:

  • The first vacuum pipes for dipoles were welded and then passed the vendor's leak test and dimensional inspection.
  • Another batch of 55 of the 1520 required bellows for the vacuum system was received for a total of 330 received to date.

Installation:

  • Quadrupole stands were installed by Bruce Co. in all available positions of the northern half of the west arc, and installation of tops and bases continued in the southern half.
  • A second wave of installation of dipole stands by Concrete Cutting Co. started in the east arc.

WBS 3

Controls:

  • Version 17 software is running in the injector but requires more testing before it's fully operational.
  • Twenty control modules have been calibrated (at a single temperature).

Fabrication:

  • Crowbar fabrication is complete.
  • We received the precision 1/2" and SMA coaxial connectors from Phoenix Logistics. They look good. The cable shop will start fabrication of the 1/2" RF feedback cables. We are now building a few test cables to confirm our tooling and process.
  • Shipped two sets of control racks to the north linac for installation in zones 17 and 18. Next week should finish the north linac.
  • Bench tested the first article of the Highland stepper motor card. It looks good. We will proceed with manufacture of the remainder of the order.

RF Power:

  • Klystron successfully passed filament life test We are nearing sign-off of agreement with Varian on changes to specs and warranty.
  • LCW pipe in north linac service building has been incorrectly welded. This is being evaluated by CEBAF and reworked with the subcontractor. It may delay availability of LCW for commissioning HPAs.

WBS 4

  • Met with box power supply vendor's representative to work out final technical details.
  • Installed another trim rack in south linac service building.
  • Completed installation of radiation monitoring conduit in tunnel.

WBS 5

No report received.

WBS 6

  • The support section for the Hall C Q1/Q2 quadrupoles was received last Friday. This is the first major structural component received for the high-momentum spectrometer (HMS). It is temporarily being stored in the north end of the EEL parking lot in preparation for Hall C's planned January 1992 move-in date. Remaining sections will be delivered during the next three months.
  • The request for proposals (RFP) for the Hall A QI quadrupole magnet was published today. Bids are due by mid-November.

WBS 7

  • Shipped a section of return transfer line to the tunnel.
  • Shipped the last supply transfer line module to the tunnel. This completes the fabrication of the supply modules.
  • Bellows valve test has completed 220,000 cycles.
  • Vacuum pumping the vacuum jacket of the return transfer line in the NW quadrant for leak testing. This test will complete the RTL in the NW quadrant.
  • CVI has completed the diode modifications and started to vacuum pump the cold box. The test schedule for the acceptance test has not been received. Possible dates are 7 through 25 October. Cooldown date still open.

WBS 8

  • Installation of crane supports in Hall A is nearing completion.
  • Continued installing the decking in preparation for placing the concrete dome on Hall B.
  • Continued post-tensioning the Hall C dome.
  • Backfilling continues in all areas when weather permits.
  • MCI continued excavating, forming, and placing ESR building footings and SB #4 footings. They also started working on SB #1 footings.

Support Services

Stockroom:

  • Withdrawal activity for the week: $18,136.59.
  • Total month expenditures: $31,976.59.
  • 297 items delivered via computer ordering methcxl.
  • Material handling effort:
    • Ten rack bases moved from Blue Crab to south linac.
    • Fifty-two cryostats moved to Fishing Point Road storage.
    • Forty-five HOM brackets moved from EEL to south linac.

External Fabrication:

  • Purchase request for beam transport drift tube supports at dipole magnet assembly written and delivery expected 1 November.
  • Drive assembly components for the first 180 degree bend have been ordered, and in-house fabricated components were completed 12 September.

Machine Shop:

  • Fourth QB magnet completed and delivered to magnet test area.
  • 1.0 meter and 1.5 meter alignment scale bars are in progress.
  • Machined, welded, and delivered four high-order sensor assemblies (kicker boxes).

Computer Center Cluster Shutdown

  • The Cluster shutdown scheduled for Tuesday, 17 September, has been rescheduled for Wednesday morning, 25 September. The entire Cluster (CEBAF/2/6,ACCELl,MICROl) will be down from approximately 8 a.m. to 10 a.m. to install a new tape drive.

CEBAF Colloquium Series

On Wednesday, 18 September, at 4 p.m. in the auditorium, Gail Hanson of Indiana State University will discuss "New Results from LEP." On Wednesday, 25 September, at 3:30 p.m. in the auditorium, Eugene Beier of the University of Pennsylvania will discuss "The Solar Neutrino Observatory."

CEBAF Science Series

The flyer for the Fall 1991 CEBAF Science Series has been distributed. The first presentation will be held at 7 p.m. on Wednesday, 25 September. Students in grades six to twelve are specifically invited, but others are also welcome. The Science Series seeks to bring students who like science together with scientists and engineers who are gifted at bringing out the fun and enjoyment of science, technology, and math. Please see the blue flyer for details.